# Semantic Structure-Mapping in LLM and Human Analogical Reasoning

### v1

@@ A.1 Regression results, Semantic Structure experiment
$smf.logit(formula=respondent\_scores\sim C(subject\_type,Treatment(reference=human))+C(quiz\_class,Treatment(reference=permuted\_questions)),data=all\_subjects\_df,).fit(maxiter=1000,method=bfgs)$
@@
$smf.logit(formula=respondent\_scores\sim C(subject\_type,Treatment(reference=human))*C(quiz\_class,Treatment(reference=permuted\_questions)),data=all\_subjects\_df,).fit(maxiter=1000,method=bfgs)$
@@
$p=chi2.sf(lik\_ratio,degfree)$, with 7 degrees of freedom.
The likelihood ratio in the above formula is calculated as follows:
$lik\_ratio=degfree*(res\_subjXclass.llf-res\_subjplusclass.llf)$.
In the above, res_subjXclass and res_subjplusclass are the regression outputs with and without interactions respectively.
For both comparisons (human subjects compared to GPT-4 and human subjects compared to Claude 3), we find a significant improvement in model fit when interactions between the subject type and experiment condition are included.A likelihood ratio test shows that including interactions between subject type and experiment condition leads to a significantly better fit of the model ($chi^{2}(7)=115.1871,p<0.001$). For the comparison between Claude 3 and human subjects, we again find a significant negative effect of the subject type being Claude 3 when interactions are not included (coef = -0.8706, z = -5.608, p < 0.001) and find that subject type - condition interactions are significant ($chi^{2}(7)=173.6511,p<0.001$). These results are consistent with the observation that the two models exhibit variable performance across conditions, and indicate that the overall performance gap to human subjects is driven by low model accuracy in certain conditions. Simple effects analysis is used below to assess the effect of subject type in particular conditions and groups thereof.
Regression outputs are shown in Figure 7 and 8.
Figure 7: Regression outputs, GPT-4 compared to human subjects in the Semantic Structure experiment.
Figure 8: Regression outputs, Claude 3 compared to human subjects in the Semantic Structure experiment.
@@ A.2 Regression results, Semantic Content experiment
Regression outputs are shown in Figure 9 and 10.
Figure 9: Regression outputs, GPT-4 compared to human subjects in the Semantic Content experiment.
Figure 10: Regression outputs, Claude 3 compared to human subjects in the Semantic Content experiment.
A.3 Further details of human performance
@@
Figure 11 shows the difference in performance between online subjects recruited through Prolific and in-person University-Name University students in the Semantic Structure experiment conditions. Prolific subjects were paid \$1.50 for the task, with Prolific taking an additional \$0.50 per subject. This equated to an approximate effective rate of \$22 per hour, well above relevant minimum wages. In-person subjects were each paid \$10 to reflect the increased time and effort cost. We expect increased performance from the in person subjects for a number of reasons. First, they are more highly remunerated. Second, there may be social pressure to perform well given the presence of a member of the research team. Third, the in-person subjects may not have the decreased attention effects likely experienced by subjects on Prolific who may complete many unrelated and potentially demotivating tasks in a day. Fourth, there is an implicit selection effect on academic performance for students at our university, which is not unrelated to the competencies involved in completing the tasks in the experiment. Indeed, we observe that accuracy increases by approximately 0.1-0.2 for the in-person subjects in all but one condition. The exception to this is the Random Finals condition, in which mean accuracy decreases slightly. However, in this condition it is not clear that a decrease in “accuracy” is objectively worse performance, because in this condition we ask for the drawing corresponding to a final unrelated term, while all previous left-hand terms within the question are related. It is thus not unreasonable to give an answer that differs from what we expect, except insofar as subjects are learning in context from previous questions in the quiz to realize that the final unrelated word should be regarded as irrelevant.
@@
Figure 11: Accuracy comparison of online subjects recruited through Prolific and in-person University-Name University students in the Semantic Structure experiment conditions.
@@
Figure 12: Human accuracy by quiz across conditions. Error bars show standard errors.
One highly-specific failure mode is present in the human data and deserves special consideration. In only the “*” grounding, participants quite often introduced separator characters into the response (either just “>” or both separator characters, “=>”). This was observed in 11 instances, thus affecting approximately 10% of the responses in that grounding scheme. This issue was not observed in any other grounding scheme. The reason for this is not entirely clear, but could be related to the short length of the groundings in this scheme (a grounding term is either 1 or 3 characters in this scheme, which is shorter than the other versions). It is possible that the short length of the grounding terms could lead subjects to perceive the separator characters as being part of the grounding term, although it is unclear why this would happen even for subjects who successfully ignore the separator characters in three prior responses (which applies to 7 out of 11 such errors).
The error rate of the human subjects by grounding type is shown in Figure 13. The error rate in the “*”, “C K E”, and “Q Z I” grounding schemes was essentially equal, with approximately half of this rate in the “c c” grounding scheme. This is not entirely intuitive, but some possible explanations of this can be offered. First, the “c c” groundings have the fewest number of distinct characters, thus limiting the option space when answering (there are two distinct characters, compared to 3 or 4 for the other groundings). Second, the specific transformations involved (capitalization and adding/removing a letter) are common operations that are encountered more frequently than, say, inserting a special character between existing characters.
@@
Some subjects who report a detailed, correct strategy nevertheless fail to attain a high accuracy, thus demonstrating that the task is not trivial even for those who are able to fully grasp what it involves. For example, one subject attains a below-average accuracy of 50% in the distracted condition despite being able to state that the task involves “Looking at other comparable entries to figure out what the answer to the last entry was (dog:puppy::cat:kitten)”, and reporting a strategy in which “I tried to find similar pairs of entries and looked at their meanings.” By contrast, another subject attains 100% accuracy in the distracted condition while responding to what the task involved with “I thought it was fun” and reporting a strategy in which “I just compared answers and tried my best to understand what they were and then tried to guess based on my interpretation of the other answers.”
@@
Figure 14: Improvement in human and LLM accuracy by question number across different conditions. Error bars show standard errors.
Further, humans and GPT-4 both improve as they see more questions over the course of a quiz. As seen in Figure 14, humans display a positive learning trend in 5 out of 8 conditions. GPT-4 displays a positive learning trend in a comparable 6 out of 8 conditions, with one of the conditions in which it does not display improvement resulting from it displaying near-perfect accuracy from start to finish (in the Only RHS condition).
A.4 Further details of LLM performance
@@
Figure 16: Dependence of GPT-4 accuracy on prompt variations. Error bars show standard errors.
@@
——— SAMPLED RESPONSE SET ———
! [Continuation omitted]
@@
——— SAMPLED RESPONSE SET ———
Table 7: Table showing an illustrative response from GPT-4 in the numerical condition of the Semantic Content experiment. Observe that the model fails to correctly relate the number of characters to the numerical property of the object, in this case the number of wheels that a car has.
Prompt * * * + * = * * * * * * * * * – * * = * * * * * * * * * – * * * =
Responses * * * * * * (first response) * * * * * * (second response) * * * * * * * (third response)

### v2

@@ A.1 Regression results, Semantic Structure experiment
smf.logit(formula=respondent_scores $\sim$
C(subject_type, Treatment(reference=human))
+ C(quiz_class, Treatment(reference=permuted_questions)),
data=all_subjects_df,
).fit(maxiter=1000, method=bfgs)
@@
smf.logit(formula=respondent_scores $\sim$
C(subject_type, Treatment(reference=human))
* C(quiz_class, Treatment(reference=permuted_questions)),
data=all_subjects_df,
).fit(maxiter=1000, method=bfgs)
@@
p=chi2.sf(lik_ratio, degfree), with 7 degrees of freedom.
where likelihood ratio in the above formula is calculated as follows:
lik_ratio = degfree * (res_subjXclass.llf - res_subjplusclass.llf).
In the equation above, res_subjXclass and res_subjplusclass are the regression outputs with and without interactions respectively.
For both comparisons (human subjects compared to GPT-4 and human subjects compared to Claude 3), we find a significant improvement in model fit when interactions between the subject type and experiment condition are included.A likelihood ratio test shows that including interactions between subject type and experiment condition leads to a significantly better fit of the model ($chi^{2}(7)=115.1871,p<0.001$). For the comparison between Claude 3 and human subjects, we again find a significant negative effect of the subject type being Claude 3 when interactions are not included (coef=-0.8706, z=-5.608, p<0.001) and find that subject type - condition interactions are significant ($chi^{2}(7)=173.6511,p<0.001$). These results are consistent with the observation that the two models exhibit variable performance across conditions, and indicate that the overall performance gap to human subjects is driven by low model match to reference in certain conditions. Simple effects analysis is used below to assess the effect of subject type in particular conditions and groups thereof.
Regression outputs are provided in the linked Github repository.
@@ A.2 Regression results, Semantic Content experiment
Regression outputs are provided in the linked Github repository.
A.3 Chain-of-thought prompting
We test the best-performing model from our experiments, Claude 3, with a chain-of-thought prompt across all experiments. We use the following prompt:
Chain-of-thought prompt: “We are conducting an experiment on general reasoning abilities. Below we will show you various words and drawings of each, after which you will need to complete the last drawing. Respond by thinking through the problem, reasoning step by step, before ending with “The answer is:” [and then the answer]. Get to the answer well within 100 words. If there is more than one question, only give the answer to the last question. When giving the answer, respond as concisely as possible with only the last drawing.”
We provide the model with a response length of 300 tokens, and grade the answer following “The answer is:” (when available, as is the case 97% of the time in the Semantic Structure experiment and 87% of the time in the Semantic Content experiment; when this phrase is not present we grade whatever is provided). The prompt components asking the model to a) get to the answer within 100 words and b) only answer the last question were added after two initial runs without these conditions resulted in the model not getting to the answer within the response window or first answering earlier questions before the live one.
The prompt yields responses that are complete and on-topic but largely wrong, with low match to reference answers across the board. This is shown in Figures 7, 8, 9, and 10 below.
Figure 7: Human and LLM reference match proportion in the Defaults and Permuted Pairs conditions. Error bars show standard errors.
Figure 8: Human and LLM reference match proportion in Only RHS, Randoms, and Random Finals conditions. Error bars show standard errors.
Figure 9: Human and LLM reference match proportion in the $2\times n$ condition followup, with Defaults condition performance for reference. Error bars show standard errors.
Figure 10: Human and model reference match proportion by condition in the Semantic Content experiment. Error bars show standard errors. Outline bars correspond to single-attribute conditions while filled bars correspond to dual-attribute conditions, with color differentiating whether numeric properties or only categorial properties are included. For example, performance in the single-attribute Numeric condition is given by the outlines surrounding peach bars while the peach bars give performance in the Numeric Dual-attribute condition. Note that “Categorial” conditions include one or two categorial properties while “Numeric” conditions include either a numeric property alone or a numeric property paired with a categorial one.
These results are surprisingly poor, and so we investigate the nature of the errors. Responses generally provide step-by-step reasoning that is related to the questions and then give an answer to the correct question, but the reasoning and answers are incorrect. An illustrative example is the following question from the Defaults condition:
Prompt: “We are conducting an experiment on general reasoning abilities. Below we will show you various words and drawings of each, after which you will need to complete the last drawing. Respond by thinking through the problem, reasoning step by step, before ending with “The answer is:” [and then the answer]. Get to the answer well within 100 words. If there is more than one question, only give the answer to the last question. When giving the answer, respond as concisely as possible with only the last drawing.
Question 1:
square => E K C
rectangle => E % K % C
circle => C K E
oval => ”
The reference answer for this question is “C % K % E”, because a rectangle is to a square as an oval is to a circle, and the transformation from square to rectangle is the insertion of percent signs.
Model responses:
Response A
“The pattern seems to be:
1. The first letter is the first letter of the shape name.
2. The middle letter is always ‘K’.
3. The last letter is the last letter of the shape name.
Applying this pattern to ‘oval’:
1. First letter of ‘oval’ is ’O’.
2. Middle letter is ‘K’.
3. Last letter of ‘oval’ is ‘L’.
The answer is: O K L”
Response B
“The pattern seems to be:
1. The first letter represents the shape (E for rectangle, C for circle).
2. The middle letter is always K.
3. The last letter is the first letter of the shape name (C for circle, E for rectangle).
Following this pattern for an oval, we would have:
1. O for oval
2. K in the middle
3. O for the last letter
The answer is: O K O”
Response C
“The pattern seems to be:
1. The first letter represents the shape (square, rectangle, circle).
2. The middle letter is always K.
3. The last letter alternates between C and E, depending on whether the shape has straight sides (C) or curved sides (E).
An oval has curved sides, so it should end with E.
The answer is: O K E”
Response D
“The pattern seems to be:
1. The first letter represents the shape (E for rectangle, C for circle).
2. The middle letter is always K.
3. The last letter is the first letter of the shape name (C for circle, E for rectangle).
Since an oval is similar to a circle, the pattern for oval should be:
The answer is: O K O”
Response E
“The pattern seems to be:
1. The first letter represents the shape (E for rectangle, C for circle).
2. The middle letter is always K.
3. The last letter is the first letter of the shape name (C for circle, E for rectangle).
Applying this pattern to an oval:
1. Oval starts with O.
2. The middle letter is K.
3. The last letter is O (first letter of “oval”).
The answer is: O K O”
It is interesting that Claude 3 performs well “intuitively” but worse when reasoning through the problems explicitly.
We test the recent Deepseek R1 “reasoning model” [56] on the same example above. Deepseek R1 is post-trained with reinforcement learning over reasoning text to improve multi-step problem solving. Interestingly and in contrast to Claude 3 with the chain-of-thought prompt, it reasons through the above prompt perfectly as shown below:
“Okay, let’s try to figure out this problem. So, there are these shapes and corresponding symbols. The first one is a square which maps to E K C. Then a rectangle is E % K % C. A circle is C K E, and the question is what the oval would be.
Hmm, I need to see the pattern here. Let’s list them out:
- Square: E K C
- Rectangle: E % K % C
- Circle: C K E
- Oval: ?
First, let’s look at the square and rectangle since they are similar shapes. The square is a four-sided figure with equal sides, and a rectangle is a four-sided figure with opposite sides equal. The symbols for square are E K C. For rectangle, it’s E % K % C. So the rectangle’s symbols have % inserted between each of the square’s symbols. So E becomes E %, then K becomes % K %, and C becomes % C? Wait, not exactly. Let me check again.
Wait, the square is E K C. The rectangle is E % K % C. So between each letter in the square’s sequence, they added a %. So E then %, then K, then %, then C. So the square’s E K C becomes E % K % C for the rectangle. That seems like inserting a % between each original character. So square to rectangle adds % between each element.
Now, circle is C K E. What would the oval be? Ovals are like stretched circles. So maybe similar to how rectangle is a stretched square. So if circle is C K E, then the oval would follow the same pattern as square to rectangle. So inserting something between the letters. But for the rectangle, they used %. Maybe for the oval, it’s a different symbol? Wait, the rectangle’s pattern was inserting % between each letter of the square. So square (E K C) becomes E % K % C. So each space between letters gets a %.
Then for circle (C K E), which is three letters, inserting % between each would be C % K % E. But wait, the rectangle’s symbols had E % K % C, which is the original square’s letters separated by %. So if we apply the same logic to the circle to get the oval, then the oval’s symbols would be C % K % E. But wait, let me check if there’s another pattern.
Alternatively, maybe the symbols are reversed for the circle. The square is E K C, and the circle is C K E. So the circle’s symbols are the reverse of the square’s. Then maybe the oval would be the reverse of the rectangle’s symbols. The rectangle is E % K % C. Reversing that would be C % K % E. So that would be the oval’s symbols. That makes sense because circle is reverse of square, so oval (which is like a stretched circle) would be the reverse of the rectangle’s symbols. So rectangle is E % K % C, reverse would be C % K % E for oval.
Yes, that seems plausible. Let me check again. Square to rectangle: inserting % between letters. Circle is reverse of square’s letters. So oval would be reverse of rectangle’s letters. Therefore, the oval’s symbols would be C % K % E. That fits the pattern. So the answer should be C % K % E.
[Output] The answer is: C % K % E”
The very good performance on this example is interesting, but we leave analysis of analogical reasoning in “reasoning models” for future work.
A.4 Further details of human performance
@@
Figure 11 shows the difference in performance between online subjects recruited through Prolific and in-person Brown University students in the Semantic Structure experiment conditions. Prolific subjects were paid \$1.50 for the task, with Prolific taking an additional \$0.50 per subject. This equated to an approximate effective rate of \$22 per hour, well above relevant minimum wages. In-person subjects were each paid \$10 to reflect the increased time and effort cost. We expect increased performance from the in person subjects for a number of reasons. First, they are more highly remunerated. Second, there may be social pressure to perform well given the presence of a member of the research team. Third, the in-person subjects may not have the decreased attention effects likely experienced by subjects on Prolific who may complete many unrelated and potentially demotivating tasks in a day. Fourth, there is an implicit selection effect on academic performance for students at our university, which is not unrelated to the competencies involved in completing the tasks in the experiment. Indeed, we observe that match to reference answers increases by approximately 0.1-0.2 for the in-person subjects in all but one condition. The exception to this is the Random Finals condition, in which mean match to reference decreases slightly. However, in this condition it is not clear that a decrease in match to reference is objectively worse performance, because in this condition we ask for the drawing corresponding to a final unrelated term, while all previous left-hand terms within the question are related. It is thus not unreasonable to give an answer that differs from what we expect, except insofar as subjects are learning in context from previous questions in the quiz to realize that the final unrelated word should be regarded as irrelevant.
@@
Figure 11: Proportion of match to reference answers comparison of online subjects recruited through Prolific and in-person Brown University students in the Semantic Structure experiment conditions.
@@
Figure 12: Human match to reference proportion by quiz across conditions. Error bars show standard errors.
One highly-specific failure mode is present in the human data and deserves special consideration. In only the “*” target domain, participants quite often introduced separator characters into the response (either just “>” or both separator characters, “=>”). This was observed in 11 instances, thus affecting approximately 10% of the responses in that target domain scheme. This issue was not observed in any other target domain scheme. The reason for this is not entirely clear, but could be related to the short length of the target domain elements in this scheme (a right hand term is either 1 or 3 characters in this scheme, which is shorter than the other versions). It is possible that the short length of the right hand terms could lead subjects to perceive the separator characters as being part of the target term, although it is unclear why this would happen even for subjects who successfully ignore the separator characters in three prior responses (which applies to 7 out of 11 such errors).
The error rate of the human subjects by target domain type is shown in Figure 13. The error rate in the “*”, “C K E”, and “Q Z I” target domain schemes was essentially equal, with approximately half of this rate in the “c c” target domain scheme. This is not entirely intuitive, but some possible explanations of this can be offered. First, the “c c” right hand terms have the fewest number of distinct characters, thus limiting the option space when answering (there are two distinct characters, compared to 3 or 4 for the other right hand terms). Second, the specific transformations involved (capitalization and adding/removing a letter) are common operations that are encountered more frequently than, say, inserting a special character between existing characters.
@@
Some subjects who report a detailed, correct strategy nevertheless fail to attain a high proportion of answers matching the reference, thus demonstrating that the task is not trivial even for those who are able to fully grasp what it involves. For example, one subject attains a below-average match to reference of 50% in the Distracted condition despite being able to state that the task involves “Looking at other comparable entries to figure out what the answer to the last entry was (dog:puppy::cat:kitten)”, and reporting a strategy in which “I tried to find similar pairs of entries and looked at their meanings.” By contrast, another subject attains 100% match to reference in the Distracted condition while responding to what the task involved with “I thought it was fun” and reporting a strategy in which “I just compared answers and tried my best to understand what they were and then tried to guess based on my interpretation of the other answers.”
@@
Figure 14: Improvement in human and LLM proportion match to reference by question number across different conditions. Error bars show standard errors.
Further, humans and GPT-4 both improve as they see more questions over the course of a quiz. As seen in Figure 14, humans display a positive learning trend in 5 out of 8 conditions. GPT-4 displays a positive learning trend in a comparable 6 out of 8 conditions, with one of the conditions in which it does not display improvement resulting from it displaying near-perfect match to reference from start to finish (in the Only RHS condition).
A.5 Further details of LLM performance
@@
Figure 16: Dependence of GPT-4 match to reference proportion on prompt variations. Error bars show standard errors.
@@
Sampled Response Set ! [Continuation omitted]
@@
Sampled Response Set * * *
Table 7: Table showing an illustrative response from GPT-4 in the Numeric condition of the Semantic Content experiment. Observe that the model fails to correctly relate the number of characters to the numerical property of the object, in this case the number of wheels that a car has.
Prompt * * * + * = * * * *